OpenStudy (calculusxy):

Given A=(1,5), B=(x,2), and C=(4,-6) and the sum of AB+BC is to be a minimum, find the value of x.

OpenStudy (sooobored):

distance formulaaaaaa \[d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\]

OpenStudy (calculusxy):

I got the distance of AB to be \(\sqrt{(x+1)^2 + 9}\)

OpenStudy (calculusxy):

and the distance of BC is \(\sqrt{(4-x)^2 + 64}\)

OpenStudy (sooobored):

i respectfully disagree with distance AB

OpenStudy (calculusxy):

What did you get?

OpenStudy (sooobored):

x-1 and 2-(-5)=7

OpenStudy (calculusxy):

I am so sorry there was a typo. I just fixed it

OpenStudy (calculusxy):

The part of the problem that I do't understand is that their sum needs to be a minimum?

OpenStudy (calculusxy):

*don't

OpenStudy (calculusxy):

Sorry I have to go. Thanks for the help :)

OpenStudy (sooobored):

well, i assume the function d(x)=AB(x)+BC(x) will look something like a positive parabola so you could use the first derivative

OpenStudy (sooobored):

and find where teh slope is equal to zero and thus the lowest point